OpenStudy (ujjwal):

I know this one involves a very simple concept but then i am totally unaware about it.. The question goes like this: The mass and volume of a body are measured and found 22.42 g and 4.7 cc respectively with error in density is: a)0.2% b)2% c)5% d)10%

OpenStudy (anonymous):

the question is incomplete

OpenStudy (anonymous):

u should be given error in mass and error in volume

OpenStudy (ujjwal):

I thought the same thing but that is the question in my book.. May be its a mistake.. Not sure though.. Well it says the answer is 2%

OpenStudy (anonymous):

no clues..

OpenStudy (anonymous):

You lot are sooo not European. Fortunately I am in the UK. The SI (ISO) open limits for error are one half the smallest measure of your equipment. So volume measured to an accuracy of .1 cc give an error of +/- .05 cc and of .01 g it is +/- 0.005 g So do the standard density m/V and then with then upper limit on mass and the lower limit of volume and see the difference. 4.770212 and 4.82258, just over +/- 1.1%, so 2%. HOWEVER: WARNING. If tou used a micrometer to measure volume then there are three measurements made, say it was a cubish shape. Then the errors are additive. .05 + .05 +.05 = 1.5. So now this would make the larger density of 4.928 g/cc. So the error would be larger at 3.3%

OpenStudy (anonymous):

lol interesting. SI units for error

OpenStudy (anonymous):

Yes if you have a rule and the smallest division on it is 1 mm then your systematic error is +/- 0.5 mm. If you have a micrometer which smallest division is 0.001" then the systematic error is +/- 0.0005"

OpenStudy (vincent-lyon.fr):

I agree with TheProf. You can also work out relative uncertainty from the formula itself. Relative uncertainty for \( \rho=\Large \frac mV\) is \( \Large \frac {\Delta \rho}{\rho}= \frac {\Delta m}{m} +\frac {\Delta V}{V} \) Here your mass is known to \(\Delta m\) = 0.01 g and your volume to \(\Delta V\) = 0.1 cm³ Uncertainty due to mass is negligible. All uncertainty is due to volume 0.1/4.7 ≈ 2%

OpenStudy (vincent-lyon.fr):

@TheProf: "You lot are sooo not European. Fortunately I am in the UK." You seem to be the first Briton I know actually showing off his/her Europeanity! No surprise you are a science teacher! Look at my dialogue with Goutham in http://openstudy.com/users/vincent-lyon.fr#/updates/4f9e75dbe4b000ae9ed2b733

OpenStudy (vincent-lyon.fr):

I will not say 'The pot calling the kettle black' anymore, but 'The Brit calling the Indian non-metric'!